OpenStudy (anonymous):

cos[sin^-1(3/5) - cos^1(3/5)]

jimthompson5910 (jim_thompson5910):

hint: let \[\Large x = \sin^{-1}\left(\frac{3}{5}\right)\] \[\Large y = \cos^{-1}\left(\frac{3}{5}\right)\]

jimthompson5910 (jim_thompson5910):

so you'll end up with cos(x-y) break this up using the identity cos(x-y) = cos(x)cos(y) + sin(x)sin(y)

OpenStudy (anonymous):

please that doesn't help i'm so lost

jimthompson5910 (jim_thompson5910):

ok let's draw a right triangle